# Emotional Dimension Control in Language Model-Based Text-to-Speech: Spanning a Broad Spectrum of Human Emotions

### v1

@@ Abstract
Current emotional text-to-speech (TTS) systems face challenges in mimicking a broad spectrum of human emotions due to the inherent complexity of emotions and limitations in emotional speech datasets and models. This paper proposes a TTS framework that facilitates control over pleasure, arousal, and dominance, and can synthesize a diversity of emotional styles without requiring any emotional speech data during TTS training. We train an emotional attribute predictor using only categorical labels from speech data, aligning with psychological research and incorporating anchored dimensionality reduction on self-supervised learning (SSL) features. The TTS framework converts text inputs into phonetic tokens via an autoregressive language model and uses pseudo-emotional dimensions to guide the parallel prediction of fine-grained acoustic details. Experiments conducted on the LibriTTS dataset demonstrate that our framework can synthesize speech with enhanced naturalness and a variety of emotional styles by effectively controlling emotional dimensions, even without the inclusion of any emotional speech during TTS training.
Index Terms:
@@
I Introduction
Imagine a world similar to the one depicted in the movie ‘Her’, where a dialogue system not only talks but connects with users emotionally [1]. Emotional text-to-speech (TTS) systems are bringing us closer to this reality [2]. Designed to generate speech from text that conveys human-like emotions, emotional TTS has seen significant improvements thanks to recent advancements in deep learning [3]. As a crucial technology for human-computer interaction, emotional TTS is paving the way for dialogue systems that can be as engaging and empathetic as the personal companion portrayed in ‘Her’ [4].
@@ I Introduction
However, current emotional TTS systems still face challenges stemming from the complexity of human emotions and limitations in both data and models [2]. Psychological studies suggest that humans can experience around 34,000 different emotions [5], yet current systems struggle to capture this full range. Previous emotional TTS research has either relied on supervising with categorical emotion labels [6, 7, 8, 9] or transferring styles from reference speech [10, 11]. Both approaches are constrained by the limited styles present in emotional speech datasets, often resulting in synthesized voices that exhibit averaged or stereotypical emotional patterns, thus providing limited control over emotion rendering [12, 13].
@@
The scarcity of emotional speech data presents significant challenges for developing emotional TTS systems [2], particularly those that utilize language models [14, 15]. With large-scale training data, recent language model (LM)-based TTS frameworks [16, 17, 18] show strong in-context learning capabilities, effectively rendering appropriate prosody for complex text inputs and exhibiting robust zero-shot speaker cloning. Unlike traditional TTS task, which primarily focuses on capturing a wide range of phonetic variations across speakers [19], emotional TTS aims to render diverse emotional styles, thus requiring a large and diverse emotional speech data with detailed annotations. Emotional TTS systems like CosyVoice [14] and EmoCtrl-TTS [15], built on state-of-the-art language models, rely on thousands of hours of emotional speech data, underscoring the complexity and cost of such data collection. However, publicly available emotional speech datasets are typically small [20] and inadequate for training a robust LM-based TTS system from scratch. A common approach is to fine-tune a pre-trained TTS model with a small emotional dataset [21], but this can lead to overfitting and limit generalization across emotions and speakers [22]. Instead of scaling emotional TTS with larger datasets to achieve state-of-the-art benchmarks, this paper integrates insights from emotion theory to enable emotional dimension control and leverages the strong in-context learning capabilities of language models to build a TTS system that spans a broad spectrum of human emotions, especially when emotional data is limited or unavailable.
In this paper, we propose an effective approach for TTS to generate diverse emotional styles by controlling a emotional dimension vector. Our proposed method does not require any emotional speech data during TTS training and demonstrates the capability to span a broad spectrum of human emotions in the synthesized voices. Our approach is particularly inspired by Russell’s studies on emotion theory [23], which suggest that the 3 independent dimensions of pleasure, arousal, and dominance adequately capture the complexity of emotional expressions and are consistent across different contexts and subjects. We leverage the definition of three dimensions denoting emotions, as reported in [23], as foundational anchors to build a model aimed at predicting these values. The TTS framework then leverages pseudo-emotional dimensions to generate fine-grained acoustic details in parallel, based on the phonetic units predicted by the autoregressive model. During inference, our proposed framework facilitate emotional dimension control to synthesize diverse emotional styles without requiring any emotional references. Our contributions are summarize as below:
  • We propose an LM-based TTS framework with emotional dimension control over pleasure, arousal, and dominance, employing a 2-stage LM training approach that effectively renders emotions without requiring any emotional training data;
  • Our experiments show that our proposed model can synthesize a wide range of emotional styles, effectively spanning a broad spectrum of human emotions;
  • We also demonstrate the effectiveness of proposed method in enhancing the naturalness and prosody consistency with the prompt speech, outperforming the baseline.
@@
Figure 1: An overview of the proposed text-to-speech (TTS) framework with emotional dimension control, consisting of: (a) Emotional Dimension (ED) Predictor Training, and (b) TTS Flow. The ED predictor is pre-trained on an emotional speech dataset to map emotional features to dimension representations via anchored dimensionality reduction. It then guides the non-autoregressive language model (LM) to predict acoustic details. ‘P’, ‘A’, and ‘D’ denote ‘Pleasure’, ‘Arousal’, and ‘Dominance’, ‘G2P’ represents ‘Grapheme-to-Phoneme’.
II Proposed Method
Following our previous research [24], we consider TTS as a conditional language modeling task, where the speech waveform is quantized into discrete tokens and is then modeled with a language model. However, training language models typically requires a large amount of data to achieve stable performance [25, 26], making it impractical as to emotional speech. Speech emotion, as a form of speech expressiveness, can be represented through continuous emotional dimensions that are learnable from expressive speech [27]. Our goal is to develop a TTS model that learns from expressive speech data during training and synthesizes speech that conveys intended emotions. The overall TTS system includes an autoregressive (AR) decoder and a non-autoregressive (NAR) decoder. The AR decoder predicts phonetic tokens from phoneme inputs and prompt audio, while the NAR decoder, using emotional dimension (ED) embeddings, predicts acoustic codecs with fine-grained details. Finally, the Encodec decoder synthesizes the speech waveform from these codecs.
II-A Predicting Emotional Dimensions from Categories
Human emotions can be represented categorically [28] or dimensionally [29]. Compared to categorical emotional approach [28] that lacks the capability to capture subtle emotional differences, the dimensional approach, such as Russell’s circumplex model, views emotions as fluid experiences across several dimensions, offering a more nuanced understanding [29]. The specific definitions [23] of 151 emotion types, as rated by 300 subjects, were reliably measured on a pleasure-arousal-dominance scale as shown in Table I. Despite those studies, most emotional TTS research still relies on a finite set of emotion labels [30, 22]. Few works have explored synthesizing mixed emotional styles by combining basic emotions [31, 32]. Current TTS systems [33, 34, 15] that control emotional dimensions rely on SER models trained with costly, labor-intensive emotional dimension annotations, and face similar biases and challenges as those encountered in SER [35, 36].
@@ II-A Predicting Emotional Dimensions from Categories
TABLE I: Some examples of the definition [23] of emotional states in terms of pleasure (‘P’), arousal (‘A’), and dominance (‘D’).
Emotion (P, A, D) Emotion (P, A, D)
Angry (-0.51, 0.59, 0.25) Elated (0.50, 0.42, 0.23)
Happy (0.81, 0.51, 0.46) Alert (0.49, 0.57, 0.45)
Sad (-0.63, -0.27, -0.33) Protected (0.60, -0.22, -0.40)
Surprise (0.40, 0.67, -0.13) Relaxed (0.68, -0.46, 0.20)
Anxious (0.01, 0.59, -0.15) Neutral (0.00, 0.00, 0.00)
II-B TTS Training with Emotional Dimension Guidance
Since speech contains both linguistic and para-linguistic information, we decouple TTS training as two stages: ‘Phoneme $\rightarrow$ Phonetic’ and ‘Phonetic $\rightarrow$ Acoustic’. The overall TTS diagram is illustrated in Figure 1(b). In the first stage, the AR model learns to map phoneme sequences to phonetic tokens. We leverage WavLM features with K-means clustering as phonetically-rich targets to train the AR model. In the second stage, the TTS focus on acoustic modeling with emotional dimension guidance, where a NAR LM is trained to recover fine-grained acoustic details from phonetic tokens predicted by AR model. Emotional dimension vectors, derived from the pre-trained ED predictor, are used to condition the NAR model, together with predicted phonetic tokens and acoustic tokens generated by Encodec. During TTS training, the AR model solely focuses on linguistic modeling, while the NAR model learns fine-grained acoustic details from the prompt audio, emphasizing the consistency of pleasure-arousal-dominance with the prompt.
II-C Emotion Rendering with Emotional Dimension Control
During inference, given input text and prompt audio along with its text transcriptions, the AR model predicts phonetic tokens representing phonetic variations over the phonemes. Subsequently, the NAR model predicts acoustic tokens based on the emotional dimension vector, the acoustic prompt, and the phoneme sequences. The emotional dimension vector can either be predicted from the prompt audio (‘ED prediction’) or manually assigned by users (‘ED control’). Finally, the encodec decoder synthesizes the speech waveform based on the predicted acoustic codecs.
III Experiments
III-A Experimental Setup
@@
The emotional dimension predictor is trained on the train set from the English subset of the ESD dataset [40], which includes around 10 hours of emotional speech data across 5 basic emotions: Neutral, Angry, Happy, Sad, and Surprise. The emotional dimension predictor reported an 81% emotion classification accuracy on the ESD test set. We train the TTS framework on LibriTTS dataset [41], which consists of around 600 hours of speech data with transcriptions at a 24kHz sampling rate from 2,456 speakers. We merge the ‘train-clean’ and ‘train-other’ subsets as the training set and evaluate TTS performance on the ‘test-clean’ subset. Note that while LibriTTS contains expressive speech, it does not include any emotion labels. There is no emotional speech included in the TTS training.
@@
  • Proposed with ED control, where ED values are assigned by users. In our experiments, ED values are directly derived from the literature [23] to simulate 5 basic emotions [28]111 5 Basic Emotions: Happy (0.81, 0.51, 0.46), Surprise (0.40, 0.67, -0.13), Angry (-0.51, 0.59, 0.25), Sad (-0.63, -0.27, -0.33), Neutral (0, 0, 0).; or redefined from the range in [23] to simulate 4 emotion pairs with subtle emotional differences2224 Emotion Pairs: Elation (1, 1, 0.5), Pleasure (1, 0, 0.5), Relaxed (0.5, -0.5, 0.2), Protected (0.5, -0.2, -0.5), Alert (0.5, 0.5, 0.5), Surprise (0.5, 0.5, -0.1), Anger (-1, 1, 1), Anxiety (-1, 1, -1), aiming to study relationships between 3 dimensions under controlled conditions;
@@
We select training speech data that is up to 10 seconds long, and all synthesized speech samples range from 3 to 10 seconds. Speech demos are publicly available online [^3] Speech Demos: https://kunzhou9646.github.io/emo-icassp25/.
III-B Implementation Details
The TTS framework consists of both AR and NAR models, both of which consist of a transformer architecture consisting of 12 layers, 16 attention heads, an embedding dimension of 1024, a feed-forward layer dimension of 4096 and a dropout rate of 0.1. The 3-D emotional dimension vector can be either extracted from the prompt audio or assigned by users, which is then projected by a linear layer of dimension 128. We use a 24kHz Encodec [^4] https://huggingface.co/facebook/encodec_24khz as the acoustic encoder to compress the speech waveform into 8 layers of discrete acoustic tokens, while the Encodec decoder is utilized to recover 8 layers of discrete acoustic tokens to the speech waveform. We obtain the phonetic tokens from the 24th layer of WavLM [^5] https://huggingface.co/microsoft/wavlm-large models and cluster them using the K-means algorithm [42]. We manually upsample the phonetic tokens from 16kHz to 24kHz to align with acoustic tokens. The WavLM model for emotional dimension prediction is fine-tuned for 100 epochs with a batch size of 64 using the Adam optimizer at a 1e-4 learning rate. Dimensionality reduction settings are a 0.1 minimum distance and 20 nearest neighbors, with a 1e-2 learning rate for point optimization.
(a) 5 basic emotions controlled by emotional dimension vector1
(b) 8 emotions controlled by emotional dimension vector2
Figure 2: Comparison of pitch, energy, and spectral flux statistics for emotions generated and controlled by the emotional dimension vector across 1,731 speech samples from the LibriTTS ‘test-clean’ set.
Figure 3: Visualization of classification probabilities for 1,731 synthesized speech from the LibriTTS ‘test-clean’, using the emotion2vec model [43], a pre-trained 9-class speech emotion recognition (SER).
III-C Objective Evaluation
We visualize the statistics of pitch, energy, and spectral flux for all synthesized results. These 3 objective metrics are widely used in the SER field, as exemplified by tools like OpenSMILE [44]. Higher pitch and energy levels often indicate ‘Happy’ or ‘Surprise’, while greater spectral flux reflects rapid changes, associated with intense emotions like ‘Excitement’. We compare 5 basic emotions, as shown in Figure 2(a). It is observed that ‘Happy’ and ‘Surprise’ exhibit higher pitch, energy, and spectral flux, while ‘Sad’ shows the lowest values among the 5 emotions. As for 4 emotion pairs shown in Figure 2(b), ‘Elation’, ‘Pleasure’, ‘Alert’, and ‘Surprise’ exhibit higher pitch, energy, and spectral flux, whereas ‘Relaxed’ and ‘Protected’ show the lowest values. These results show that our model can synthesize distinguishable emotions solely by adjusting the emotional dimensions, aligning with the expected emotional characterizations.
We further employ a publicly available pre-trained speech emotion recognizer (SER), Emotion2vec [^6] https://huggingface.co/emotion2vec/emotion2vec_base, to evaluate all synthesized results. Emotion2vec [43], trained on 9 categories, is used to visualize the classification probabilities for 5 basic emotions, as shown in Figure 3. For 3 out of 5 intended emotions, the corresponding classification probabilities are the highest, demonstrating that our synthesized emotions can be perceptually recognized by the SER.
(a) Mean opinion score (MOS) to evaluate speech naturalness
(b) XAB Preference to evaluate speech naturalness and prosody similarity with the prompt speech
Figure 4: Subjective evaluation results assess the zero-shot performance on the TTS task. In the proposed framework, the emotional dimension (ED) is predicted based on the prompt speech (‘Proposed w/ ED Prediction’).
TABLE II: Comparative Mean Opinion Score (CMOS) for assessing emotional intelligibility of 4 synthesized emotions (‘Angry’, ‘Happy’, ‘Sad’, ‘Surprise’) compared to synthesized neutral speech.
Emotion Angry Happy Sad Surprise
CMOS 0.64 $\pm$ 0.33 0.71 $\pm$ 0.25 0.15 $\pm$ 0.32 0.92 $\pm$ 0.38

### v2

# Zhou Zhang Zhao Wang Pan Ng Zhang Ni Ma Nguyen Yip Ma Alibaba GroupSingapore RochesterUnited States of America

@@ Abstract
Current emotional text-to-speech (TTS) systems face challenges in conveying the full spectrum of human emotions, largely due to the inherent complexity of human emotions and the limited range of emotional labels in existing speech datasets. To address these limitations, this paper introduces a TTS framework that provides flexible user control over three emotional dimensions—pleasure, arousal, and dominance—enabling the synthesis of a diverse array of emotional styles. The framework leverages an emotional dimension predictor, trained solely on categorical labels from speech data and grounded in earlier psychological research, which is seamlessly integrated into a language model-based TTS system. Experimental results demonstrate that the proposed framework effectively learns emotional styles from expressive speech, eliminating the need for explicit emotion labels during TTS training, while enhancing the naturalness and diversity of synthesized emotional speech.
keywords:
@@
1 Introduction
Emotional text-to-speech (TTS) aims to generate speech from text that conveys human-like emotions [1]. Recent advances in deep learning have significantly enhanced its capabilities. As a pivotal technology for human-computer interaction, emotional TTS facilitates the development of dialogue systems that are both engaging and empathetic [2].
@@ 1 Introduction
However, current emotional TTS systems face challenges stemming from the complexity of human emotions and the limited range of emotional labels available in existing databases [2]. Psychological research suggests that humans can experience around 34,000 distinct emotions [3], yet current systems struggle to capture this full spectrum. Previous emotional TTS studies have primarily relied on two approaches: supervision using categorical emotion labels [4, 5] and style transfer from reference speech [6, 7]. Both approaches are constrained by the limited diversity of emotional styles present in available datasets, often resulting in synthesized speech with averaged or stereotypical emotional expressions, which restricts fine-grained control over emotion rendering [8, 9].
@@
Recent language model (LM)-based TTS frameworks [10, 11] show strong in-context learning capabilities. With sufficient training data, these models effectively render appropriate prosody for complex text inputs and exhibit robust zero-shot speaker cloning capabilities. Unlike traditional TTS task, which focus primarily on capturing a wide range of phonetic variations across speakers [12], emotional TTS emphasizes rendering diverse emotional styles, thus requiring large and varied emotional speech datasets with detailed emotion annotations. Emotional TTS systems like CosyVoice [13] and EmoCtrl-TTS [14], built on state-of-the-art language models, require thousands of hours of emotional speech data, highlighting the complexity and cost of data collection. However, publicly available emotional speech datasets are typically small [15], making them inadequate for training robust LM-based TTS system from scratch. A common strategy is to fine-tune a pre-trained TTS model using small emotional datasets [16], but this approach risks overfitting and limits generalization across emotions and speakers [17]. To address these limitations, we propose a novel framework that synthesizes more nuanced and subtle emotions not present in the dataset by controlling continuous emotional dimensions. This method offers greater flexibility for emotion rendering and leverages the in-context learning capabilities of language models, enabling the generation of a wide spectrum of human emotions, even with limited or no emotion-labeled data.
In this paper, we propose an effective approach for LM-based TTS systems to generate diverse emotional styles by controlling an emotional dimension vector. Our method draws inspiration from Russell’s emotion theory [18], which identifies three independent dimensions—pleasure, arousal, and dominance—as sufficient to capture the complexity of emotional expressions across various contexts and individuals. Using these dimensions as foundational anchors, we develop a model that predicts their values. The pre-trained dimension predictor is then integrated into a LM-based TTS system to assist in next-token prediction, alongside the speaker embedding. During inference, our framework enables direct control over emotional dimensions, allowing the synthesis of diverse emotional styles without requiring emotional reference inputs. Our contributions are summarize as below:
  • We introduce a novel approach for controlling emotional dimensions in language model-based TTS systems, enabling fine-grained emotional speech synthesis;
  • Our approach learns emotional styles directly from expressive speech, eliminating the need for explicit emotion labels during TTS training;
  • During inference, the model allows users to control and explore a broad spectrum of human emotions by adjusting the emotional dimension vector. Grounded in established psychological theory, this approach ensures that the synthesized emotional styles align with the full range of human emotions;
  • We further demonstrate that our method enhances naturalness in zero-shot speaker cloning tasks, consistently outperforming the baseline system;
@@
Figure 1: An overview of the proposed text-to-speech (TTS) framework with emotional dimension control, consisting of: (a) Emotional Dimension (ED) Predictor Training, and (b) Text-to-Speech Flow. The ED predictor is pre-trained on an emotional speech dataset to map emotional features to dimension representations via anchored dimensionality reduction. It then guides the autoregressive language model (LM) to predict acoustic details. ‘P’, ‘A’, and ‘D’ denote ‘Pleasure’, ‘Arousal’, and ‘Dominance’.
2 Proposed Method
2.1 Problem Formulation
We model TTS as a conditional language modeling task, where the speech waveform is quantized into discrete tokens and is then modeled with a language model [10, 13, 12]. However, training language models typically requires a large amount of data to achieve stable performance [19, 20], making it impractical as to emotional speech. Speech emotion, as a form of speech expressiveness, can be represented through continuous emotional dimensions that are learnable from expressive speech [21]. Our goal is to develop a TTS model that learns from expressive speech data during training and synthesizes speech that conveys intended emotions. Similar with that of CosyVoice [13], the overall TTS system includes an autoregressive LM decoder and a flow-matching module. The autoregressive LM decoder predicts acoustic tokens from text inputs and prompt speech using emotional dimension (ED) embeddings. Subsequently, a flow-matching module predicts the Mel-spectrum from the discrete acoustic tokens. Finally, the HiFi-GAN vocoder synthesizes the speech waveform from the Mel-spectrum.
2.2 Predicting Emotional Dimensions from Categories
Human emotions can be represented categorically [22] or dimensionally [23]. Compared to categorical emotional approach [22] that lacks the capability to capture subtle emotional differences, the dimensional approach, such as Russell’s circumplex model, views emotions as fluid experiences across several dimensions, offering a more nuanced understanding [23]. The specific definitions [18] of 151 emotion types, as rated by 300 subjects, were reliably measured on a pleasure-arousal-dominance scale as shown in Table 1. Despite those studies, most emotional TTS research still relies on a finite set of emotion labels [24, 17]. Few works have explored synthesizing mixed emotional styles by combining basic emotions [25, 26]. Current TTS systems [27, 28, 14] that control emotional dimensions rely on SER models trained with costly, labor-intensive emotional dimension annotations, and face similar biases and challenges as those encountered in SER [29, 30].
@@ 2.2 Predicting Emotional Dimensions from Categories
Table 1: Examples of the definition [18] of emotional states in terms of pleasure (‘P’), arousal (‘A’), and dominance (‘D’).
Emotion P A D Emotion P A D
Angry $-0.51$ $0.59$ $0.25$ Excited $0.62$ $0.75$ $0.38$
Happy $0.81$ $0.51$ $0.46$ Alert $0.49$ $0.57$ $0.45$
Sad $-0.63$ $-0.27$ $-0.33$ Protected $0.60$ $-0.22$ $-0.40$
Surprise $0.40$ $0.67$ $-0.13$ Relaxed $0.68$ $-0.46$ $0.20$
Anxious $0.01$ $0.59$ $-0.15$ Neutral $0.00$ $0.00$ $0.00$
2.3 TTS Training with Emotional Dimension Guidance
Following CosyVoice [13], we decompose the TTS process into two stages: ‘text-to-token’ and ‘token-to-speech’. The overall TTS diagram is illustrated in Figure 1(b). Given prompt speech, emotional dimension vectors are extracted using the pre-trained ED predictor (Figure 1(a)), while the speaker embedding is obtained from a pre-trained voiceprint model [^1] https://github.com/alibaba-damo-academy/3D-Speaker/tree/main/egs/3dspeaker/sv-cam++. The input text is first processed through a Grapheme-to-Phoneme tokenizer and then mapped to a shared semantic space with speech tokens using a text encoder. These extracted features collectively serve as conditioning inputs for the speech tokenizer, which produces speech token sequences.
We then frame the TTS task as an auto-regressive speech token generation task using a language model (LM). During LM training, we employ a teacher-forcing scheme to optimize LM predictions, where the left-shifted sequence serves as the input, and the original sequence acts as the expected output. We compute the cross-entropy loss between the predicted and ground-truth speech tokens, ensuring effective learning of token dependencies and accurate reconstruction of speech sequences.
2.4  Emotion Rendering with Emotional Dimension Control
During inference, the emotional dimension vector can be obtained in two ways: either be predicted from the prompt speech (‘ED prediction’) or manually assigned by users (‘ED control’), allowing for flexible emotion customization. Given the input text and prompt speech, the language model (LM) autoregressively predicts a sequence of speech tokens, which encode both linguistic and prosodic information. These tokens are then processed by a pre-trained optimal-transport flow-matching model, which learns the underlying distribution of Mel spectrograms. The flow-matching model generates high-fidelity Mel spectrogram samples that are conditioned on the predicted speech tokens, ensuring smooth and natural prosody. Finally, the synthesized Mel spectrograms are passed through a HiFi-GAN vocoder, which converts them into a high-quality speech waveform. Our model is expected to generate expressive and natural speech while preserving the desired emotional characteristics, whether inferred from the prompt speech or explicitly controlled by the user.
Figure 2: Statistical analysis of pitch, energy, and spectral flux for nine synthesized emotions, obtained by modulating the emotional dimension vector within the proposed framework using ED control (’ED Control’).
3 Experiments
Table 2: Mean Opinion Score (‘MOS’) of naturalness for the zero-shot TTS task. In the proposed system, emotional dimension values are predicted from the prompt speech (‘ED Prediction’).
Ground Truth CosyVoice [13] Proposed
4.80 $\pm$ 0.08 4.36 $\pm$ 0.13 4.54 $\pm$ 0.18
Figure 3: Mean Opinion Score for assessing emotional intelligibility (‘E-MOS’). In the proposed system, ED values were predicted based on the emotional reference (‘ED Prediction’).
3.1 Experimental Setup
@@
The emotional dimension predictor is trained on the train set from the English subset of the ESD dataset [34], which includes around 10 hours of emotional speech data across 5 basic emotions: Neutral, Angry, Happy, Sad, and Surprise. The emotional dimension predictor reported an 81% emotion classification accuracy on the ESD test set. We train the TTS framework on LibriTTS dataset [35], which consists of around 600 hours of speech data with transcriptions at a 24kHz sampling rate from 2,456 speakers. We merge the ‘train-clean’ and ‘train-other’ subsets as the training set and evaluate TTS performance on the ‘test-clean’ subset. Note that while LibriTTS only contains expressive speech, it does not include any explicit emotion labels.
@@
  • Proposed with ED control, where ED values are directly derived from the literature [18] to simulate different emotions, as shown in Table 1;
@@
All synthesized speech samples range from 3 to 10 seconds. Speech demos are publicly available online [^2] Speech Demos: https://demos46.github.io/is25/.
3.2 Implementation Details
The backbone TTS framework was implemented based on the official release of CosyVoice [13], which consists of a speech tokenizer, a text encoder, a text-to-token LM and a flow-matching module. The speech tokenizer employs a ESPNet Conformer ASR model along with a vector quantizer of a codebook size of 4096 after the first 6 encoder layers. The text encoder consists of 6 transform layers with an 8-head and 512-d attention layer. The LM contains 12 transform layers with an 8-head and 512-d attention layer. An optimal-transport conditional flow matching model (OT-CFM) [36] is employed to predict the Mel spectrograms. A HiFi-GAN [^3] https://github.com/jik876/hifi-gan vocoder recovers the speech signal from the predicted Mel spectrogram.
The emotional dimension predictor consists of a WavLM features extractor followed by a linear layer, and an anchored dimensionality reduction module. The WavLM model is fine-tuned for 100 epochs with a batch size of 64 using the Adam optimizer at a 1e-4 learning rate. Dimensionality reduction settings are a 0.1 minimum distance and 20 nearest neighbors, with a 1e-2 learning rate for point optimization.
3.3 Objective Evaluation
We compute the statistics of pitch, energy and spectral flux for all synthesized speech, as shown in Figure 2. These acoustic features play a crucial role in characterizing emotional expressiveness. Higher pitch and energy levels are commonly associated with emotions such as ‘Excitement’, where increased vocal effort and higher frequency components convey heightened arousal. Similarly, greater spectral flux, which reflects rapid spectral variations, is indicative of dynamic speech patterns, often linked to intense emotions like ‘Surprise’, which exhibit abrupt prosodic changes.
From Figure 2, we observe that emotions such as ‘Excited’, ‘Surprise’, ‘Happy’ and ‘Alert’ exhibit relatively higher values across all three acoustic features, suggesting that these emotions are characterized by greater vocal dynamism and expressiveness. In contrast, emotions like ‘Sad’, ‘Angry’, ‘Neutral’ and ‘Protected’ tend to have lower values, indicating more subdued, steady, or restrained speech patterns. For instance, ‘Sad’ is typically marked by lower pitch and energy, corresponding to a softer, more monotonous voice, while ‘Neutral’ maintains moderate values across all features, aligning with a balanced, unexpressive tone. These findings provide quantitative evidence that our model successfully synthesizes distinguishable emotions solely by adjusting the emotional dimensions. The alignment between our results and established emotional speech patterns suggests that the model effectively captures the expected emotional characterizations.